### v3

# A Theory of Quantum Jumps and Its Applications to Lyman and Balmer Series

Abstract
A theory of quantum jumps is developed by using a new asymmetric equation, which is complementary to the Schrödinger equation. The new equation displays Bohr’s rules for quantum jumps, and its solutions demonstrate that once a quantum jump takes place at a random time, then its evolution is continuous and coherent. The temporal solutions are used to determine time-scales of the quantum jumps, which are very short but finite. The spatial solutions are used to find the radial probability density of finding the electron after its jump. The obtained results are applied to the Lyman and Balmer series absorption, and their experimental verification is discussed.
1 Introduction
The randomness and abruptness of quantum jumps between two discrete energy levels in an atom was the main feature of quantum theory originally formulated by Bohr [1] in 1913. His main argument was that electrons change their allowed energy levels by absorbing or emitting photons of very specific energies, and that the quantum jumps corresponding to absorption and emission are instantaneous. This notion of random and instantaneous quantum jumps became one of the pillars foundation of the Copenhagen interpretation of quantum mechanics (QM) (e.g., [2-5]). However, Schrödinger [6] argued against such instantaneous quantum jumps by pointing out that no process in the real world happens instantenously or in zero time.
The first direct observations of quantum jumps in individual atoms exposed to electromagnetic (EM) fields took place in 1986 and were reported in [7-9]. The collected data showed that the atoms flipped between a state that allowed them to emit a photon to a state that did not, and that the atoms remain in one state, or the other, for periods ranging from a few tenths of a second to a few seconds before jumping again. Other indepedent experiments confirmed the results and demonstrated that the jumps were random and abrupt (e.g., [10-12], and references therein). A different view has emerged from the experimental results obtained by Minev et al [13], who confirmed that quantum jumps come at random times, but once they come, the evolution of each completed jump is a continuous and coherent physical process that unfolds in a finite time. It was suggested that a quantum trajectories theory could account for the experimental data [14-16].
The theoretical approach described in this paper is different from any previous attempt to account for quantum jumps, including the quantum trajectories theory. The new theory of quantum jumps developed herein is based on a new asymmetric equation (NAE) that was derived from the irreducible representations (irreps) of the extended Galilean group [17]. Since the same irreps can be used to derive the Schrödinger equation (SE), and since there are other similarities between the equations, the SE and NAE are considered to be complementary to each other. The main physical implication of this complementarity is that quantum phenomena, like quantum jumps or quantum measurements, are not described by the SE, but they may be represented by the NAE.
The NAE naturally describes interactions of quantum systems with environments that are subjected to monitoring by measuring devices. Specifically, the NAE can directly describe interactions of a single quantum particle with a measuring apparatus, or the process of absorption or emission of electromagnetic (EM) waves by quantum systems. Thus, while the SE is used to formulate unitary quantum theories, the NAE can also be used to formulate non-unitary theories because of its second-order derivatives in time (see Section 2, for details). Random and instantaneous quantum jumps violate the unitary evolution (e.g., [2-13]), thus, an attempt herein to describe them by using the NAE. Similarly, the quantum measurement problem is a non-unitary evolution of a quantum system during its measurement (e.g., [18-21]) and, therefore, the NAE was used to solve it [22].
It was recently demonstrated that a non-unitary theory of quantum measurements can be formulated using the NAE, and that its predictions are consistent with Born’s position measurement postulates [22]. Moreover, since the origin and nature of dark matter remains unknown (e.g., [23-26]), and since all previous attempts to formulate a theory of dark matter (DM) based on the SE (e.g., [27,28]) failed (e.g., 29,30]), a theory of DM using the NAE was developed [22], and applied to construct models of galactic cold DM halos [31,32], whose theoretical predictions are proposed to be verifed by specific astronomical observations [33]. The NAE was also used to describe propagation of classical (acoustic) waves in isothermal and stratified media [34].
The main aim of this paper is to develop a theory of quantum jumps using the NAE. The theory is applied to a hydrogen atom and used to describe absorption Lyman and Balmer series. It is shown that Bohr’s rule for quantum jumps (e.g., [3-5]) emerges naturally from the NAE, and that the NAE can be solved analytically by separation of variables. To preserve the coherence of the wavefunction, the solutions are restricted to the EM frequencies corresponding directly to the jumps; in other words, measurements that destroy the coherence are not accounted for by the solutions. The temporal solutions of the NAE show that time-scales for the quantum jumps are very short but finite. The spatial solutions are used to obtain the radial probability densities, which are normal distributions with the location of their centers in agreement with Bohr’s rule for quantum jumps. Experimental verification of the obtained results is also discussed.
The paper is organized as follows: the basic equations resulting from the Galilean group are given and discussed in Section 2; the governing equation and its solutions are presented in Section 3; the solutions are applied to the Lyman and Balmer absorption processes in Section 4; experimental verification of the obtained results is discussed in Section 5; and Conclusions are given in Section 7.
2 Schrödinger and new asymmetric equations
In Galilean relativity, space and time are separated and described by different metrics, which are invariant with respect to rotations, translations and boosts that form the Galilean group (e.g., [35,36]). The irreducible representations (irreps) of the Galilean group were originally obtained and classified by Bergmann [37], who demonstrated that only scalar and spinor irreps are physical in Galilean space and time. The spinor irreps were used by Lévy-Leblond [38,39] to obtain his spinor wave equation that describes elementary particles with spins in nonrelativistic QM. In this paper, only the scalar irreps are considered.
In the extended Galilean group (e.g., [38,39]), the space and time translations are combined into one subgroup, $T(3+1)$, whose properties are the same as the $T(3+1)$ subgroup of the Poincaré group [35,36]. Since $T(3+1)$ is an invariant subgroup, its irreps are the irreps of the entire extended Galilean group, and they can be used to derive equations of nonrelativistic and relativistic QM [40]. According to Wigner [41], an elementary particle is uniquely identified if its wavefunction transforms as one of the irreps of the extended Galilean group. For a scalar wavefunction $\phi(t,\mathbf{x})$, the Wigner condition is given by the following eigenvalue equations $i\partial_{t}\phi=\omega\phi$ and $-i\nabla\phi=\mathbf{k}\phi$, where $\partial_{t}=\partial/\partial t$, and $\omega$ and $\mathbf{k}$ are the eigenvalues that label the irreps [42,43].
Using the eigenvalue equations and the de Broglie relationship (e.g., [3-5]), the following two asymmetric equations are derived [17,32]:
$\left[i{{\partial}\over{\partial t}}+\frac{\hbar}{2m}\nabla^{2}\right]\phi_{S}(t,\mathbf{x})=0\ ,$ (1)
which is the Schrödinger equation (e.g., [3-5]), and
$\left[\frac{i}{\omega}{{\partial^{2}}\over{\partial t^{2}}}+\frac{\hbar}{2m}{\mathbf{k}}\cdot\nabla\right]\phi_{A}(t,\mathbf{x})=0\ ,$ (2)
which is called the new asymmetric equation or NAE [31-33]. Since both equations are derived from the same irreps, they may be considered as complementary to each other. However, the main difference between them is that the SE is the first-order in its time derivative, thus allows for unitary theories, while the NAE can also be used to construct non-unitary theories because of its second-order time derivative. In addition, the NAE requires specifying the eigenvalues $\omega$ and $\mathbf{k}$ that are explicitly present in the equation.
There are processes in nonrelativistic QM, such as the quantum measurement problem and quantum jumps, that are considered to be non-unitary [3-5]. In these processes, the external input/output of EM radiation plays important role, and the NAE directly allows accounting for this radiation by specifying its $\omega$ and $\mathbf{k}$. A possible solution of the quantum measurement problem by using the NAE is presented in [22]. In this paper, the NAE is used to formulate a theory of quantum jumps.
3 Basic equations and their solutions
3.1 The governing equation
To apply the NAE to quantum jumps, it is required that the labels of the irreps $\omega$ and $\mathbf{k}$ are specified. Let $\omega=\omega_{o}$, $E_{o}=\hbar\omega_{o}$, and ${\mathbf{k}}={\mathbf{k_{o}}}$, where $\omega_{o}$ and ${\mathbf{k_{o}}}$ represent frequency and wavevector of EM radiation absorbed or emitted by a hydrogen atom. Then, Eq. (2) can be written in the following form
$\left[\frac{i\hbar}{\omega_{o}}\left({{\partial^{2}}\over{\partial t^{2}}}\right)+\frac{\hbar^{2}}{2m}({\mathbf{k}_{o}}\cdot\nabla)+V(r)\right]\phi_{A}(t,\mathbf{r})=0\ ,$ (3)
where $V(r)=-e^{2}/(4\pi\epsilon r)$ is the Coulomb potential.
To determine $\omega_{o}$ and ${\mathbf{k_{o}}}$, one may use
$\Delta E=R_{E}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right)\ ,$ (4)
where $R_{E}=\hbar^{2}/2ma_{B}^{2}$ is the Rydberg constant, $a_{B}=(4\pi\epsilon_{o}\hbar^{2})/me^{2}$ is the Bohr radius, and $n_{i}$ and $n_{f}$ are the prinicipal quantum numbers corresponding to the initial and final states, respectively. For absorption, $n_{f}>n_{i}$ and $\Delta E>0$; however, for emission, $n_{f}<n_{i}$ and $\Delta E<0$. Taking $\Delta E=\hbar\omega_{o}$, one obtains
$\omega_{o}=\frac{R_{E}}{\hbar}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right)\ ,$ (5)
and
${\mathbf{k_{o}}}=k_{o}{\mathbf{\hat{k}}_{o}}=\frac{\omega_{o}}{c}{\mathbf{\hat{k}}_{o}}=\frac{R_{E}}{c\hbar}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right){\mathbf{\hat{k}}_{o}}\ .$ (6)
Then, Eq. (3) can be written as
$\left[\frac{i\hbar}{\omega_{o}}\left({{\partial^{2}}\over{\partial t^{2}}}\right)+\left(\frac{\lambda_{C}}{4\pi}\right)(\hbar\omega_{o})({\mathbf{\hat{k}}_{o}}\cdot\nabla)-\left(\frac{\hbar^{2}}{ma_{B}}\right)\frac{1}{r}\right]\phi_{A}(t,\mathbf{r})=0\ ,$ (7)
which is the governing equation for the presented theory, with $\lambda_{C}=h/mc$ denoting the Compton wavelength.
3.2 Temporal and spatial solutions
To solve the governing equation, the spherical variables are separated into the temporal and spatial (radial only) components, $\phi_{A}(t,\mathbf{r})=\chi(t)\ \eta(\mathbf{r})$, and the separation constant $-\mu^{2}=E_{n}=-(\hbar^{2}/2m)(1/na_{B})^{2}$. Then, Eq. (7) becomes
${{d^{2}\chi}\over{dt^{2}}}+i\left(\frac{\hbar\omega_{o}}{mn^{2}a_{B}^{2}}\right)\chi=0\ ,$ (8)
and
${{d\eta}\over{dr}}+\frac{2}{(\mathbf{\hat{k}_{o}}\cdot\mathbf{\hat{r}})k_{o}a_{B}}\left(\frac{1}{n^{2}a_{B}}-\frac{1}{r}\right)\eta=0\ ,$ (9)
where $n$ is the principle quantum number, and $n=n_{f}$ for absorption and $n=n_{i}$ for emission; since the solutions also depend on $n$, one gets $\chi(t)=\chi_{n}(t)$ and $\eta(r)=\eta_{n}(r)$.
Taking $i=(1/\sqrt{2}+i/\sqrt{2})^{2}$, Eq. (8) can be integrated giving solutions
$\chi_{n}(t)=C_{\pm}\exp{\left[\pm\ i\left(\frac{1}{\sqrt{2}}+\frac{i}{\sqrt{2}}\right)\Omega_{n}t\right]}\ ,$ (10)
where $C_{\pm}$ are the integration constants corresponding to the $\pm$ solutions, and the characteristic frequency $\Omega_{n}$ is given by
$\Omega_{n}=\frac{1}{na_{B}}\sqrt{\frac{\hbar\omega_{o}}{m}}\ .$ (11)
Both solutions with $C_{+}$ and $C_{-}$ are physical and they correspond to $t\rightarrow+\infty$ and $t\rightarrow-\infty$, respectively. In the following, only the solution with $C_{+}$ is considered because quantum jumps occur when $t>0$. The real part of the solution is
${\cal{R}\it e}[\chi_{n}(t)]=C_{+}\cos{(\Omega_{n}t)}\exp{(-\Omega_{n}t)}\ ,$ (12)
and its physical meaning and applications to a hydrogen atom are presented and discussed in Setion 3.
Before solving Eq. (9), it must be noted that in spherical symmetry $\mathbf{\hat{r}}$ can always be aligned with $\mathbf{\hat{k}_{o}}$, which means that $\mathbf{\hat{k}_{o}}\cdot\mathbf{\hat{r}}=1$. It is also convenient to transform Eq. (9) to the new independent variable $r_{a}=r/a_{B}$, which gives
$\frac{d\eta_{n}}{dr_{a}}+\beta_{B}\left(\frac{1}{n^{2}}-\frac{1}{r_{a}}\right)\eta_{n}=0\ ,$ (13)
where
$\beta_{B}=8\pi\left(\frac{a_{B}}{\lambda_{c}}\right)\left(\frac{n_{f}^{2}n_{i}^{2}}{n_{f}^{2}-n_{i}^{2}}\right)\ .$ (14)
According to Eq. (13), $\eta_{n}(r)$ reaches its maximum value when $r_{a}=n^{2}$ or $r=n^{2}a_{B}$, which is Bohr’s rule for quantum jumps (e.g., [3-5]). In other words, Eq. (13) directly displays the Bohr rule that gives the most probable radius $r_{a}$, where the maximum value of $\eta_{n}(r)$ occurs.
The solution to Eq. (13) is
$\eta_{n}(r)=\eta_{o,n}\ r_{a}^{\beta_{o}}\ \exp{\left[-\left(\frac{\beta_{o}}{n^{2}}\right)r_{a}\right]}\ ,$ (15)
where the integration constant $\eta_{o,n}$ is obtained from the following normalization condition
[…]
For $\eta_{n}(r)$ given by Eq. (15), the integral can be evaluated and the result is
$\eta_{o,n}^{2}(\beta_{B})=\frac{1}{4\pi a_{B}^{3}\Gamma(2\beta_{B}+3)}\left(\frac{2\beta_{B}}{n^{2}}\right)^{2\beta_{B}+3}\ ,$ (17)
which is valid if ${\cal{R}\it e}(2\beta_{o})>0$ and ${\cal{R}\it e}(4\beta_{o}+2)>-1$; note that both conditions are satisfied in the theory presented herein.
3.3 Radial probability density
Having obtained the normalized spatial wavefunction $\eta_{n}(r)$, the radial probability density in a spherical shell volume element can be evaluated, and it is given by
$dP_{n}(r_{a})=4\pi a_{B}^{3}r_{a}^{2}|\eta_{n}(r_{a})|^{2}dr_{a}\ ,$ (18)
or, after using Eqs (15) and (17), one obtains
${\cal{P}}_{n}(r_{a})\equiv\frac{dP_{n}(r_{a})}{dr_{a}}=\frac{1}{\Gamma(2\beta_{B}+3)}\left(\frac{2\beta_{B}}{n^{2}}\right)^{2\beta_{ap}+3}\ r_{a}^{2(\beta_{ap}+1)}\ e^{-(2\beta_{ap}/n^{2})r_{a}}\ ,$ (19)
[…]
3.4 Physical meaning of the solutions
Having obtained the temporal and spatial solutions to the governing equation, the physical meaning of these solutions is now discussed. According to Eq. (12), the temporal component $\chi_{n}(t)$ of the wavefunction $\phi_{A}(t,\mathbf{r})$ decays exponentialy in time, and the rate of this exponential decay depends on the value of the characteristic frequency $\Omega_{n}$ given by Eq. (11). Let the time $T_{n}=2\pi/\Omega_{n}$ be the duration time of a single quantum jump; the results presented below show that the duration time is very short but finite. By calculating this time for a considered quantum jump, and knowing the distance travelled by the electron during the jump, the electron speed can be determined for each jump.
The spatial solution and its normalization factor given by Eqs (15) and (17), respectively, were used to obtain the radial probability density for any quantum jump (see Eq. 19). The theory predicts that after an electron makes a quantum jump, the probablility density of finding it is very narrow and centered at $r=n^{2}a_{B}$, which is consistent with Bohr’s rule for quantum jumps. Experimental results [6-8] demonstrated that the electron remains in the excited state for periods ranging from a few tenths of a second to a few seconds before jumping again, and returns to its original state. The developed theory describes such physical situations, and its solutions preserve the coherence of the wavefunction. To achieve it, the solutions are restricted only to the EM frequencies corresponding directly to the quantum jumps (see Sec. 4). In other words, the solutions do not account for measurements because they destroy the coherence of the wavefunction.
In the following, the obtained theoretical results are applied to a hydrogen atom and its Lyman and Balmer series absorption. The theory also describes emission of EM radiation when the initial, $n_{i}$, and final, $n_{f}$, are properly defined for such transitions.
4 Application to hydrogen atom
4.1 Lyman series absorption
The developed theory is now applied to a hydrogen atom by considering absorption of EM radiation that causes the electron’s transitions from its initial, $n_{i}=1$, to its final, $n_{f}=2$, $3$ or $4$ energy levels. For this Lyman series absorption, the frequencies $\omega_{o}$ and $\Omega_{n}$ are calculated (see Eqs 5 and 11) by using
$\omega_{o}=2.065\cdot 10^{16}\left(\frac{n_{f}^{2}-n_{i}^{2}}{n_{f}^{2}n_{i}^{2}}\right)\ ,$ (20)
and
$\Omega_{n}=2.03\cdot 10^{8}\ \frac{\sqrt{\omega_{o}}}{n}\ ,$ (21)
where $n=n_{f}$. The characteristic frequency $\Omega_{n}$ can be used to estimate the duration time of quantum jumps $T_{n}=1/\Omega_{n}$. Knowing the distance traveled by the electron and the duration time, the electron’s speed $v_{e}=(n_{f}^{2}-n_{i}^{2})a_{B}/T_{n}$ can also be calculated.
Lyman series $n=n_{f}$ $\omega_{o}$ [$s^{-1}$] $\Omega_{n}$ [$s^{-1}$] $T_{n}$ [s] $v_{e}$ [$m\ s^{-1}$]
Ly-$\alpha$ 2 $1.549\cdot 10^{16}$ $1.265\cdot 10^{16}$ $5.0\cdot 10^{-16}$ $3.2\cdot 10^{5}$
Ly-$\beta$ 3 $1.836\cdot 10^{16}$ $9.171\cdot 10^{15}$ $6.9\cdot 10^{-16}$ $6.2\cdot 10^{5}$
Ly-$\gamma$ 4 $1.936\cdot 10^{16}$ $7.061\cdot 10^{15}$ $8.9\cdot 10^{-16}$ $8.9\cdot 10^{5}$
Table 1. Theoretical predictions of the characteristic frequency $\Omega_{n}$, the duration time $T_{n}$, and the speed of electron $v_{e}$ during quantum jumps corresponding to the Lyman series resulting from absorption of EM radiation with frequency $\omega_{o}$.
The obtained results are presented in Table 1. It shows that the characteristic frequency $\Omega_{n}$ of quantum jumps is of the order of $(10^{15}-10^{16})$ $s^{-1}$ for the considered transitions in the Lyman series absorption; the frequency decreases for the quantum jumps to higher values of $n_{f}$. Then, the corresponding duration time $T_{n}=2\pi/\Omega_{n}$ is of the order of $10^{-16}$ $s$, which is very short but yet finite.
Since the distance the electron travels during one specific jump is known, and since the duration time of this jump is also estimated, the electron’s speed is calculated. For the three transitions considered for the Lyman series absorption, the speed is of the order of $10^{5}$ $m/s$ as shown in Table 1; its value increases for the transition to higher orbitals.
To calculate the radial probability density given by Eq. (19), it is required to determine $\beta_{B}$, which is defined by Eq. (14) and written as
$\beta_{B}=548.2\left(\frac{n_{f}^{2}n_{i}^{2}}{n_{f}^{2}-n_{i}^{2}}\right)\ .$ (22)
Figure 1: The normalized wavefunction, $\eta_{n}(r_{a})$, in units $a_{B}^{-3/2}$, is plotted versus the ratio of radius to the Bohr radius, $r/a_{B}$, for the Lyman series absorption from $n_{i}=1$ to $n_{f}=2$, $3$ and $4$.
Figure 2: The radial probability density ${\cal{P}}_{n}(r_{a})$ in a spherical shell volume element corresponding to each transition is plotted versus the ratio of radius to the Bohr radius, $r/a_{B}$. The plotted probability densities are for the Lyman series absorption from $n_{i}=1$ to $n_{f}=2$, $3$ and $4$.
Having obtained $\beta_{B}$, the wavefunction $\eta_{n}(r_{a})$ given by Eq. (15) is calculated for the transitions from $n_{1}=1$ to $n_{f}=2$, $n_{f}=3$ and $n_{f}=4$. The wavefunctions corresponding to these transitions are normalized (see Eq. 17) with its units being $a_{B}^{-3/2}$ (e.g., [4]), and plotted in Fig. 1. The presented results demonstrate that the maxima of the wavefunctions are centered at $r_{a}$ = $4$, $9$ and $16$, respectively, which is in agreement with Bohr’s rules for quantum jumps. Moreover, the wavefunctions are symmetric about their corresponding maxima, and the shapes of the wavefunctions change from very narrow for the lowest $n_{f}$ to much wider for larger $n_{f}$.
The computed wavefunctions are then used to calculate the radial probability density ${\cal{P}}_{n}(r_{a})$ in a spherical shell volume element corresponding to each transition by using Eq. (19). The obtained results are ploted in Fig. 2. These are the probability densities of finding the electron at its new location after the EM radiation required for the jump was absorbed. The theory predicts that the resulting radial probablility densities are centered at $r_{a}$ = $4$, $9$ and $16$, respectively, as postulated by Bohr in his rules for quantum jumps (e.g., [4]). The shapes of these probability curves range from very narrow for the lowest $n_{f}$ to much wider for higer values of $n_{f}$. Comparison of the results of Fig. 2 to those in Fig. 1 shows the effect of $r_{a}^{2}$ on the plotted ${\cal{P}}_{n}(r_{a})$.
The theory predicts the characteristic frequency ($\Omega_{n}\sim 10^{16}$ $s^{-1}$), the time-scale ($T_{n}\sim 10^{-16}s$) for the considered jumps, which are used to estimate the electron’s speed during the jumps; according to Table 1 the speeds are $v_{e}\sim 10^{5}$ $m/s$. Moreover, the theory also gives the radial probability density of finding the electron in the excited state. As the experiments [7-9] demonstrated, the electron spends from a few tenths of a second to a few seconds in the excited state before it jumps back to its original orbital.
4.2 Balmer series absorption
The theory is now applied to the Balmer series absorption for the transitions from $n_{i}=2$ to $n_{f}=3$, $4$ and $5$,which correspond to the $H\alpha$, $H\beta$ and $H\gamma$ Balmer lines, respectively. The obtained characteristic frequency, $\Omega_{n}$, the time-scale, $T_{n}$, and the electron’s speed, $v_{e}$, are given in Table 2, which shows that $\Omega_{n}\sim 10^{15}$ $s^{-1}$, $T_{n}\sim 10^{-15}$ $s^{-1}$, and $v_{e}\sim 10^{5}$ $m/s$. Comparison of these results to those obtained for the Lyman series absorption (see Table 1) indicates that while there is one order of magnitude difference in $\Omega_{n}$ and $T_{n}$, the electron’s speed is of the same order for the all quantum jumps.
Balmer series $n=n_{f}$ $\omega_{o}$ [$s^{-1}$] $\Omega_{n}$ [$s^{-1}$] $T_{n}$ [s] $v_{e}$ [$m\ s^{-1}$]
[…]
Table 2. Theoretical predictions of the characteristic frequency $\Omega_{n}$, the duration time $T_{n}$, and the speed of electrons $v_{e}$ during quantum jumps corresponding to the Balmer series resulting from absorption of EM radiation with frequency $\omega_{o}$.
Figure 3: The normalized wavefunction, $\eta_{n}(r_{a})$, in units $a_{B}^{-3/2}$, is plotted versus the ratio of radius to the Bohr radius, $r/a_{B}$, for the Balmer series absorption from $n_{i}=2$ to $n_{f}=3$, $4$ and $5$.
Figure 4: The radial probability density ${\cal{P}}_{n}(r_{a})$ in a spherical shell volume element corresponding to each transition is plotted versus the ratio of radius to the Bohr radius, $r/a_{B}$. The plotted probability densities are for the Balmer series absorption from $n_{i}=2$ to $n_{f}=3$, $4$ and $5$.
The wavefunction $\eta_{n}(r_{a})$ for the transitions from $n_{1}=2$ to $n_{f}=3$, $n_{f}=4$ and $n_{f}=5$ is computed using Eq. (15) and plotted in Fig. 3; the presented wavefunctions are normalized (see Eq. 17) and their units are $a_{B}^{-3/2}$ (e.g., [4]). The location of the maximum of each plotted wavefuction is centered at $r_{a}$ = $9$, $16$ and $25$, respectively, which is consistent with Bohr’s rules for quantum jumps. Just as for the wavefunctions for the Lyman series absorption (see Fig. 1), the wavefunctions for the Balmer series absorption are also symmetric about their maxima, and their shapes change from very narrow for the lowest $n_{f}$ to much wider for larger $n_{f}$. These may be the characteristic features of all the series absorptions in a hydrogen atom.
Having obtained the normalized wavefunctions for the Balmer series absorption, Eq. (19) is used to calculate the radial probability density ${\cal{P}}_{n}(r_{a})$ in a spherical shell volume element corresponding to each transition, and the results are plotted in Fig. 4. As shown, the probability densities demonstrate that the most probable location of the electron after its absorption of the required EM radiation is $r_{a}$ = $9$, $16$ and $25$, respectively, which is in agreement with Bohr’s rules for quantum jumps (e.g., [4]). The shapes of these probability curves change from very narrow for the lowest $n_{f}$ to much wider for higer values of $n_{f}$, similar as the probability densities for the Lyman series absorption shown in Fig. 2. Moreover, comparing Figs 4 and 3, the effect of $r_{a}^{2}$ on the plotted ${\cal{P}}_{n}(r_{a})$ is clearly seen.
5 Experimental verification
According to the results presented in Tables 1 and 2, the duration time, $T_{n}$, for the electron’s transitions corresponding to the Lyman and Balmer series absorptions is finite and of the order of $10^{-16}$ $s$ and $10^{-15}$ $s$, respectively. The fact that the duration time is finite, not instantenous as originally postulated (e.g., [1-5]), is consistent with the recent experimental results [13]. However, the predicted values of $T_{n}$ for the hydrogen atom cannot be directly compared to the times of quantum jumps measured by [13]. The reason is that the measurements in [13] were performed on the so-called artificial atoms, which are much larger in size than atoms, but they have discrete energy leveles. The times $T_{n}$ given in Tables 1 and 2 are theoretical predictions based on the developed theory and they still remain to be verfied experimentally.
The measurements performed by three different teams [7-9] demonstrated that individual atoms, after absorbing EM radiation, stay in their excited states for periods ranging from a few tenths of a second to a few seconds before jumping again. Several other indepedent experiments [10-12] confirmed this range. The theoretical results presented in this paper give the radial probablility densities in spherical shell volume elements for the Lyman and Balmer series absorptions, and these probability densities may last for time periods determined by the experiments. It is proposed that the predicted radial probability densities for a single quantum jump can be verified by using a quantum microscope similar to that designed by researchers [44], who used it to measure the orbital structure of Stark states in an excited hydrogen atom. There are other methods to measure quantum phenomena [45-48], but some of them may not be suitable to observe single orbitals in a hydrogen atom [47].
6 Conclusions
In this paper, a new asymmetric equation, which is complementary to the Schrödinger equation, is used to develop a theory of quantum jumps. The main advantage of this new theory is that it explicitly displays the Bohr rules for quantum jumps and it allows for the theory to be non-unitary. The solutions to the new equation are used to determine the time-scales of quantum jumps, and to calculate the radial probability density of finding the electron after its quantum jump. The obtained solutions are applied to the Lyman and Balmer series.
The theoretical results obtained in this paper are qualitatively in agreement with the experimental results given by Minev et al. [13], who showed that quantum jumps come at random times, but once they come, the evolution of each completed jump is a continuous and coherent physical process that takes place in a finite time. The specific values of the duration times of the considered quantum jumps in the Lyman and Balmer series are predicted to be in the range of $10^{-16}-10^{-15}$ $s$. Since the measurements in [13] were performed on the so-called artificial atoms, which are much larger than atoms, the theoretically predicted times cannot be directly compared to the experimental results; thus, they they still remain to be verfied experimentally.
In accordance with the experiment results [6-9], the theoretically predicted radial probability densities are considered to last for periods ranging from a few tenths of a second to a few seconds; the experiments showed that after such times, the electron returns to its original orbital. Thus, it is proposed that the presented radial probability densities be observed experimentally by using a quantum microscope similar to that designed in [44], or other methods [45-48] suitable to observe single orbitals in a hydrogen atom.
Acknowledgment: The author thanks Dora Musielak for valuable comments on the earlier version of this manuscript.